OpenStudy (anonymous):

How do you solve this differential equation: (x^3)y''+(x^2)y'-xy=x?

zepdrix (zepdrix):

\[\large x^3y''+x^2y'-xy=x\]Divide through by \(\large x\),\[\large x^2y''+xy'-y=1\]The complimentary solution \(\large y_c\) is given by,\[\large x^2y''+xy'-y=0\]I think this part shouldn't be too difficult, it's a Cauchy Euler Equation. Do you know how to solve those? :o

OpenStudy (anonymous):

No, we haven't gone over them yet.

zepdrix (zepdrix):

Hmm :o

OpenStudy (anonymous):

That's ok. I was just running through some practice problems and happened upon this one, though nothing like it was taught in the lesson. So it's not critical :) Thanks for your help, though. I'm going to be logging out pretty soon.

zepdrix (zepdrix):

ah i see c: